a 1.00 g sample of a metal x (that is known to form x+2 ions) was added to a 0.100cm3 of 0.500mol dm3 h2so4. after all the metal head reacted, the remaining acid required 33.40 cm3l of 0.500mol dm3 naoh solution for neutralization.calculate the molar mass of the metal and identify the element.
1
Expert's answer
2020-04-06T12:01:58-0400

The reaction given can be written as following:

X + H2SO4 --> H2 + XSO4

The number of moles of H2SO4:

n(H2SO4) = C(H2SO4) × V(H2SO4) = 0.500 mol/dm3 × 100 cm3 = 0.500 mol/L × 0.100 L = 0.05 mol

The number of moles of NaOH use to neutralize H2SO4 equals:

n(NaOH) = C(NaOH) × V(NaOH) = 0.500 mol/dm3 × 33.40 cm3 = 0.500 mol/L × 33.40 × 10-3 L = 0.0167 mol

The neutralization reaction is as following:

H2SO4 + 2NaOH --> Na2SO4 + 2H2O

As a result, 0.0167 mol of NaOH neutralized 0.0167 / 2 = 0.00835 mol of H2SO4.

From here, 0.05 mol - 0.00835 mol = 0.04165 mol of H2SO4 reacted with a metal X. As a result, 0.04165 mol of metal X was present.

Molecular weight of metal X equals:

Mr(X) = m(X) / n(X) = 1.00 g / 0.04165 mol = 24 g/mol

Metal X is magnesium (Mg)


Answer: Magnesium (Mg), 24 g/mol